You are given the following cost data:

q TFC TVC
0 25 0
1 25 7
2 25 12
3 25 18
4 25 25
5 25 34
6 25 46
7 25 62
8 25 88

a) If the price of output is $15, how many units of output will this firm produce?

b) What is the total revenue? What is the total cost?

c) Will the firm operate or shut down in the short run? in the long run? Briefly explain your answers.
Answer #1

TC = TFC + TVC

MC = Change in TC / Change in Q

Q TFC TVC TC MC
0 25 0 25
1 25 7 32 7
2 25 12 37 5
3 25 18 43 6
4 25 25 50 7
5 25 34 59 9
6 25 46 71 12
7 25 62 87 16
8 25 88 113 26

a. When P = 15, Setting P = MC for profit maximization,

We see from table that P is not equal to MC, so we select output level, where MC is just below P

So output level = 6

b. TR when output is 6 = 6 * 15 = 90

TC when output is 6 = 71

c. The firm will operate in the short run as it is earning an economic profit, In long run, more firms will enter the market and firm will produce where economic profit is equal to zero.